# DeepVigor+: Scalable and Accurate Semi-Analytical Fault Resilience Analysis for Deep Neural Networks

### v1

@@ V-E Impact of Input Data on the Quality of Results
To obtain the VFs in DeepVigor+, we considered one batch of 100 data. Nonetheless, to investigate the impact of data on the quality of analysis results, we repeat the experiments for different batches of data with the size of 100 input data and derive DNNs’ total MVF. Fig. 10 illustrates the obtained total MVF for DNNs over 8 different batches of data. As observed, the variation between the total MVF for each DNN is negligible for different batches of data, demonstrating that analyzing DNNs’ resilience with one batch of data provides confident results.
VGG-11VGG-16ResNet-18-CMobileNet-V2ResNet-18-IResNet-34
1111222233334444555566667777888811111.41.41.41.41.81.81.81.82.22.22.22.22.82.82.82.83333Batch numberMVF (%)
@@
Figure 10: Total MVF variation over different batches of data for all DNNs.
V-F DeepVigor+ Applications and Considerations
As shown, DeepVigor+ achieves a fast, scalable and accurate resilience analysis for emerging DNNs. The analysis provided by DeepVigor+ is not limited to fault resilience assessment, but it can be exploited for designing fault-tolerant and resilient DNNs. Identifying more vulnerable channels and layers can lead to cost-effective selective fault mitigation techniques as well as a comparative investigation between different architectures of DNNs. Moreover, it enables design space exploration to identify more resilient DNNs against faults.
The other output of DeepVigor+ is VVR representing the values that a fault should affect neuron/weight to misclassify DNN’s golden results. These values can be used for obtaining VFs and identifying more vulnerable components as well as for fault detection at inference and identifying bit-level resilience analysis since they are represented in the power of 2 values and can be mapped down to bits.
@@ V-F DeepVigor+ Applications and Considerations
Yet, there are some constraints in this method to be considered and extended in future research:
@@
  • It is assumed that the parameters within the layers of DNNs under analysis are unifiedly distributed among channels and their places are not resorted after training. In such cases, a higher channel sampling ratio is needed to obtain accurate VF results.
@@
  • DeepVigor+ supports a single-bit fault model in input activations and weights of convolutional layers and obtained VF corresponds to this fault model. For multi-bit fault models, the corresponding error propagation should be applied.
@@
  • The error propagation analysis presented in this paper is based on 32-bit floating point data representation. However, the same concept can be extended and applied to fixed-point and integer data representations for QNNs resilience analysis, as in [49].
VI Conclusion
@@
This paper addresses one of the major challenges in fault resilience analysis for DNNs in the literature. It introduces DeepVigor+, the first semi-analytical scalable alternative method to fault injection, quantifying emerging DNN’s resilience accurately in a short time. DeepVigor+ is facilitated by optimal fault propagation modeling in DNNs accompanied by stratified sampling tackling the scalability problem for resilience analysis for DNNs. This open-source method unleashes a fast resilience assessment, enabling fine-grain evaluation and design space exploration for various fault-tolerant and cost-effective designs for DNNs.
@@
The results in the paper indicate that DeepVigor+ derives vulnerability factors for layers and the entire model of DNNs with less than 1% error, with 14.9 up to 26.9 times fewer simulations than the best-known state-of-the-art statistical FI. It is shown that DNN’s Model Vulnerability Factor can be obtained within minutes by analyzing their activations and weights. DeepVigor+ is presented as an open-source tool for researchers and engineers to enable them to exploit it for DNNs’ fault resilience assessment and enhancement.

### v2

@@ V-E Impact of Input Data on the Quality of Results
To obtain the VFs in DeepVigor+, this paper has considered one batch of $100$ input data. To investigate the impact of data on the quality of analysis results, we repeat the experiments using different batches of data with $100$ input data and derive the total MVF of the CNNs. Fig. 10 illustrates the obtained total MVF for CNNs over $8$ different batches of data. As observed, the variation between the total MVF for each CNN is negligible across different batches of data, demonstrating that analyzing CNNs’ resilience with a single batch of data provides confident results.
To evaluate the sensitivity of DeepVigor+ to the number of input samples, we conducted experiments using varying batch sizes ranging from $20$ to $200$ across all CNN models. The results revealed a negligible variation in the computed MVF, with differences consistently below $0.05\%$ across all configurations. These findings indicate that the resilience analysis performed by DeepVigor+ is largely insensitive to the batch size and the specific set of input images used, demonstrating the robustness and stability of the method.
VGG-11VGG-16ResNet-18-CMobileNet-V2ResNet-18-IResNet-3400.50.5111.51.5222.52.533Total MVF (%)
@@
Figure 10: Total MVF variation over different batches of data for all CNNs, when batch-size = $100$.
VI Discussion
VI-A Constraints and Considerations
@@
As shown, DeepVigor+ achieves a fast, scalable, and accurate resilience analysis for CNNs. Yet, there are some constraints in this method to be considered and extended in future research:
@@
  • It is assumed that the parameters within the layers of CNNs under analysis are uniformly distributed among channels and their positions are not rearranged after training. In opposite cases, a higher channel sampling ratio is needed to obtain accurate VF results.
@@
  • DeepVigor+ supports the single-bit fault model in input activations and weights of convolutional layers, and the obtained VFs correspond to this fault model. For multi-bit fault models, the corresponding error propagation should be applied.
@@
  • The error propagation analysis presented in this paper is based on 32-bit floating-point data representation. However, the same concept can be extended and applied to fixed-point and integer data representations for QNNs resilience analysis, as in [51].
  • DeepVigor+ is presented for CNNs on the image classification application. Yet, the presented theories and techniques for fault propagation are valid for other types of DL models and applications, such as RNNs and Transformers in object detection, language processing, etc. To apply DeepVigor+ to other DL models and applications, the error propagation mechanism and failure analysis should be extended based on their unique operations.
VI-B Applications, Impacts, and Future Directions
The proposed DeepVigor+ framework introduces a scalable and semi-analytical method for evaluating the fault resilience of CNNs. It offers a powerful alternative to traditional FI techniques. DeepVigor+ enables rapid and accurate reliability assessments of deep and complex CNNs at the early stages of design. This capability is particularly impactful for emerging AI and DL systems, where fast prototyping and early validation are critical for both research and deployment.
The analytical theories of DeepVigor+ open promising directions for extending reliability analysis beyond CNNs. Emerging DL models such as Transformers are increasingly gaining traction in various domains, including natural language processing, vision, and multimodal learning. Therefore, there is a growing need for efficient reliability assessment tools. The theory developed in DeepVigor+, particularly the optimal fault propagation analysis, can be generalized to these architectures with minimal overhead, primarily by integrating their unique operations into the framework. This adaptability makes DeepVigor+ a foundational tool for reliability analysis in the next generation of DL systems.
DeepVigor+ can play a significant role in quantifying the reliability of a DL system. DeepVigor+ provides the probabilistic metrics (i.e., VF) for failure due to faults at the DL model. Therefore, by combining platform-specific fault models and deployment constraints, DeepVigor+ could be integrated into design-time tools for co-optimizing efficiency and reliability. By developing cross-layer analysis, from algorithm to hardware implementation, it is possible to quantify how hardware architecture and mappings affect the reliability of the DL system. This comprehensive view facilitates the path toward certifying DL systems under emerging safety and reliability standards, such as those used in automotive or healthcare applications.
Beyond reliability analysis, DeepVigor+ can be exploited for designing fault-tolerant CNNs. Identifying the most vulnerable channels and layers by DeepVigor+ can lead to selective error detection and correction for the most vulnerable components in a DNN accelerator. While redundancy is quite costly for DNN accelerators, selective mechanisms can enable cost-effective fault tolerance.
Looking ahead, future research directions include extending DeepVigor+ to handle stochastic and temporal DL models, integrating it with real-world deployment environments (e.g., edge or neuromorphic hardware), and incorporating reliability metrics into Neural Architecture Search (NAS). These directions will further consolidate DeepVigor+ as a necessary component in building robust, safe, and certifiable AI systems.
VII Conclusion
@@
This paper addresses one of the major challenges in fault resilience analysis for CNNs. It introduces DeepVigor+, a semi-analytical, scalable alternative method to fault injection, quantifying deep CNNs’ resilience accurately in a short time. DeepVigor+ is facilitated by a theoretical fault propagation modeling in CNNs, accompanied by stratified sampling, tackling the scalability problem for their resilience analysis. This open-source method enables fast resilience assessment, providing fine-grain evaluation and design space exploration for various fault-tolerant and cost-effective designs.
@@
The results in the paper indicate that DeepVigor+ derives vulnerability factors for layers and the entire model of CNNs with $14.9$ up to $26.9$ times fewer simulations than the best-known state-of-the-art statistical FI under similar less than $1\%$ error constraints. It is shown that CNN Model Vulnerability Factors can be obtained within minutes by analyzing their activations and weights. DeepVigor+ is also made public as an open-source tool for researchers and engineers to enable them to exploit it for CNNs’ fault resilience assessment and enhancement.